OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = x + 8 y = x + 3 y = −x + 8 y = −x + 3 Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −x − 2 y = 2x − 2 y = −x − 7 Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = x − 3 y = x + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

These Also Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (−6, 3). y = x − 1 y = x + 3 y = −x − 1 y = −x + 3 Explain, in complete sentences, the relationship between the equations of perpendicular lines.

OpenStudy (anonymous):

parallel will have the same slope, and perpendicular will have the negative reciprocal